OpenStudy (yanni):

The ratio of men to women working for a company 2 to 3 . If there are 125 employees total, how many women work for the company?

OpenStudy (shane_b):

If the ratio of men to women is 2:3, the there are 3 women for every 5 people. You can simply divide 125 by 5 and multiply that times 3 to get the number of women.

OpenStudy (shane_b):

That's one method anyway.

OpenStudy (yanni):

75?

OpenStudy (shane_b):

Yes

OpenStudy (lgbasallote):

the ratio is 2:3 and the total is 125 therefore let 2x = number of men 3x = number of women x is the ratio thingy 2x + 3x = 125 5x = 125 x = 5 now to find the number of women simply do 3x and sub in x = 5 women = 3(5)

OpenStudy (shane_b):

I just knew @lgbasallote would write it all out :)

OpenStudy (lgbasallote):

haha

OpenStudy (yanni):

so there are 15 women

OpenStudy (lgbasallote):

i dont have the enthusiasm to write it in separate

OpenStudy (lgbasallote):

no x is actually 25

OpenStudy (shane_b):

Right...the answer is still 75 no matter which way you do it.

OpenStudy (yanni):

wait...is the answer 75, 15, or 25

OpenStudy (lgbasallote):

i was just trying to see if someone will notice ;)

OpenStudy (lgbasallote):

it's 75

OpenStudy (shane_b):

75...@lgbasallote meant to write 3(25)

OpenStudy (lgbasallote):

x = 25 3x = 3(25) = 75

OpenStudy (yanni):

oh kk thx

OpenStudy (shane_b):

I didn't notice it until yanni did :P